# Sound Speed Resonance of Gravitational Waves in Gauss–Bonnet-coupled inflation

### v1

@@ I Introduction
In this paper, we put forward a specific class of models dynamically realizing the SSR mechanism. We show how SSR is obtained in case of inflaton coupled to a Gauss–Bonnet (GB) term. GB-coupled theories were extensively studied in Refs. [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] as a viable candidate for inflation. Here, we will consider a simple set of models which can be naturally implemented in the inflationary $\alpha$-attractor paradigm inspired by Refs. [28, 29, 30] (although our mechanism is not limited to $\alpha$-attractors). We will also examine the case of double inflation with the first inflaton coupled to the GB term. In this case, we will demonstrate how characteristic GW signals by SSRs are obtained. These can be tested in future pulsar-timing radio-astronomy experiments or space-based interferometers.
@@ III Scalar modes
In order to avoid the singularities in the evolution of perturbations during oscillations of the inflaton, one can use the Mukhanov variable [39]
@@
as noticed in Refs. [35, 36]. Here $\delta\phi$ and $\psi$ are gauge-invariant inflaton and metric perturbations, respectively. Following Ref. [37], we work in longitudinal gauge where the metric takes the form,
$ds^{2}=-(1+2\Phi)dt^{2}+a^{2}(1-2\Psi)dx^{2}~{}.$ (18)
We then obtain a (linearized) coupled system of equations for $\Psi$ and $\delta\phi$,
@@ V Benchmark model
with some constants $\lambda$ and $n>0$. During early inflation, when $\phi\ll 1$, the GB coupling is suppressed, while towards the end of slow-roll it can be activated for appropriately chosen $\lambda$.
By choosing $\alpha=n=1$, in Figure 1 we show the numerical background solution $\phi$ (top-left), GB slow-roll parameter $\omega$ ($\epsilon$ is shown in the inset plot) (top-right), sound speed of scalar modes (bottom-left), and sound speed of tensor modes (bottom-right). The plots show normalized time near the end of inflation, $\epsilon(t_{\rm end})=1$, and $\tilde{\lambda}\equiv\lambda M^{2}$. The plots of $C_{s}$ and $C_{t}$ show their qualitatively different behaviour: $C_{s}^{2}$ has a single dip (with barely visible features afterwards), while $C_{t}^{2}$ oscillates with gradually decaying amplitude. This difference can be explained by noticing in (11) and (36) that the time-dependent part of $C_{s}^{2}$ is at least quadratic in $\omega$, i.e. $C_{s}^{2}=1+{\cal O}(\omega^{2})$, while for $C_{t}^{2}$ we have $C_{t}^{2}=1+{\cal O}(\omega)$. Since the magnitude of $\omega$ is quickly reduced after initial spike (see Figure 1), the oscillations of $C_{s}^{2}$ are suppressed by $\omega^{2}$, while the oscillations of $C_{t}^{2}$ are suppressed by $\omega$.
@@ VI Two-field inflation and observable SGWB
The resulting numerical plots are similar to those shown in Figure 1, with the exception that $t_{\rm end}$ now indicates the end of the first stage, and $\epsilon$ starts to decrease after some oscillations, due to the resumption of slow-roll. Figure 2 shows resonant amplification of tensor perturbations for the case $n=1$ (first row) and $n=2$ (second row), with $\alpha=1$ and $\gamma=0.01$ in both cases (we set the initial value $\chi_{0}=9$ which results in $\sim 22$ efolds of $\chi$-driven second phase of inflation). The first column shows GW power spectrum in terms of the density fraction $h^{2}\Omega_{\rm GW}$, and the second column shows the respective squared sound speeds around the end of the first phase of inflation. The inflaton mass parameter is adjusted around $M\sim 1.3\times 10^{-5}$ by CMB normalization of the scalar power spectrum, $P_{\zeta}(k_{\rm CMB})\approx 2.1\times 10^{-9}$. The squared sound speeds of tensor and scalar modes are positive in both cases of Figure 2. We find taht the scalar perturbations do not experience any amplification in our models and do not contribute to either PBH abundance or secondary GWs.
@@
In principle, by testing different scalar potentials and GB functions, one can obtain a larger amplification (in fact, slightly increasing $\gamma$ in (42) can also lead to somewhat larger peaks, as will be shown below). For example, we find that for
@@ VII Conclusions
First, we have demonstrated a time-variation of the sound speeds of scalar and tensor modes driven by the Lagrangian interactions, by focusing on the regime of damped oscillations of the inflaton around its minimum. These oscillations lead to damped oscillations of the tensor mode sound speed, while the variations of the scalar mode sound speed are suppressed by comparison. Then, we showed how the GB-induced SSR can generate a large stochastic background of GWs if the oscillatory phase is followed by more e-folds, as can be realized in two-field double inflation scenario. Notably, we obtained several examples of GW signals which can be tested in future experiments in radio astronomy and space-based interferometers  [^2] An alternative GW phenomenology of (single) GB-coupled inflation, in the frontier of ultra-high-frequencies (UHF) far beyond LIGO, was recently studied in Refs. [51, 52]. Moreover UHF GWs can also have an important impact in radio astronomy from stochastic resonant graviton-photon transitions with cosmological magnetic fields [53, 54].. This represents an important step towards the embedding of the SSR phenomenon, for GWs and PBHs co-genesis, in a theory of inflation based on a Lagrangian principle and fundamental symmetries. Several results show how inflationary $\alpha$-attractors can be obtained from models based on superconformal groups, supergravity and string theory [55, 56, 57], while Gauss–Bonnet portals can be generated by string theory corrections [58]. These encourage the future exploration of SSR in string-inspired inflation scenarios.

### v2

@@ I Introduction
In this paper, we put forward a specific class of models dynamically realizing the SSR mechanism. We show how SSR is obtained in case of inflaton coupled to a Gauss–Bonnet (GB) term. GB-coupled theories were extensively studied in Refs. [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] as a viable candidate for inflation or dark energy. Here, we will consider a simple set of models which can be naturally implemented in the inflationary $\alpha$-attractor paradigm inspired by Refs. [35, 36, 37] (although our mechanism is not limited to $\alpha$-attractors). We will also examine the case of double inflation with the first inflaton coupled to the GB term. In this case, we will demonstrate how characteristic GW signals by SSRs are obtained. These can be tested in future pulsar-timing radio-astronomy experiments or space-based interferometers.
Unless otherwise stated, we work in Planck units, $M_{P}=1/\sqrt{8\pi G}=1$.
@@ III Scalar modes
In order to avoid the singularities in the evolution of perturbations during oscillations of the inflaton, one can use the gauge-invariant Mukhanov variable [46]
@@
as noticed in Refs. [42, 43]. Here $\delta\phi$ and $\Psi$ are gauge-invariant inflaton and spatial metric perturbation (see Eq. (18)), respectively. Following Ref. [44], we work in longitudinal gauge where the metric takes the form,
$ds^{2}=-(1+2\Phi)dt^{2}+a^{2}(1-2\Psi)dx^{2}~{},$ (18)
while the gauge-invariant perturbation $\delta\phi$ from (17) coincides with the usual inflaton perturbation, so we denote both as $\delta\phi$. We then obtain a (linearized) coupled system of equations for $\Psi$ and $\delta\phi$,
@@ V Benchmark model
with some constants $\lambda$  [^2] Upon restoring the Planck mass, we write the Lagrangian as ${\cal L}\sim\frac{M_{P}^{2}}{2}(R+\xi R_{\rm GB}^{2})$, so that the parameter $\lambda$ has the dimension [mass]-2. and $n>0$. During early inflation, when $\phi\ll 1$, the GB coupling is suppressed, while towards the end of slow-roll it can be activated for appropriately chosen $\lambda$.
By choosing $\alpha=n=1$, in Figure 1 we show the numerical background solution $\phi$ (top-left), GB slow-roll parameter $\omega$ ($\epsilon$ is shown in the inset plot) (top-right), sound speed of scalar modes (bottom-left), and sound speed of tensor modes (bottom-right). The plots show normalized time near the end of inflation, $\epsilon(t_{\rm end})=1$, and $\tilde{\lambda}\equiv\lambda M^{2}$ ($\tilde{\lambda}$ is dimensionless, since in our convention $[\lambda]=[{\rm mass}]^{-2}$). The plots of $C_{s}$ and $C_{t}$ show their qualitatively different behaviour: $C_{s}^{2}$ has a single dip (with barely visible features afterwards), while $C_{t}^{2}$ oscillates with gradually decaying amplitude. This difference can be explained by noticing in (11) and (36) that the time-dependent part of $C_{s}^{2}$ is at least quadratic in $\omega$, i.e. $C_{s}^{2}=1+{\cal O}(\omega^{2})$, while for $C_{t}^{2}$ we have $C_{t}^{2}=1+{\cal O}(\omega)$. Since the magnitude of $\omega$ is quickly reduced after initial spike (see Figure 1), the oscillations of $C_{s}^{2}$ are suppressed by $\omega^{2}$, while the oscillations of $C_{t}^{2}$ are suppressed by $\omega$.
@@ VI Two-field inflation and observable SGWB
The resulting numerical plots for background fields are similar to those shown in Figure 1, with the exception that $t_{\rm end}$ now indicates the end of the first stage, and $\epsilon$ starts to decrease after some oscillations, due to the resumption of slow-roll.
In order to derive the gravitational wave power spectrum, we numerically solve Eq. (35), where the slow-roll parameters and the scale factor are given by the (numerical) background solution of the two-field model (42). We start numerical integration of (35) deep inside the horizon, e.g. at the time $t_{i}$ when $C_{t}k=100aH$, and stop the integration when the mode is well outside the horizon, e.g. at the time $t_{f}$ when $C_{t}k=aH/100$. The exception is when $t_{i}$ and $t_{f}$ fall in the period of transient slow-roll violation (when $\phi$ starts oscillating around its local minimum), in which case we move $t_{i}$ well before the slow-roll violation happens (such that the mode is in the Bunch–Davies vacuum), and $t_{f}$ well after the slow-roll regime is re-established (to allow the solution to stabilize after the oscillatory phase). The value of the solution $y$ is then taken at the end of numerical integration ($t_{f}$).
The resulting GW density fraction is shown in Figure 2, where resonant amplification of tensor modes can be seen for the cases $n=1$ (first row) and $n=2$ (second row), with $\alpha=1$ and $\gamma=0.01$ in both cases (we set the initial value $\chi_{0}=9$ which results in $\sim 22$ efolds of $\chi$-driven second phase of inflation). The first column shows GW power spectrum in terms of the density fraction $h^{2}\Omega_{\rm GW}$, and the second column shows the respective squared sound speeds around the end of the first phase of inflation. The inflaton mass parameter is adjusted around $M\sim 1.3\times 10^{-5}$ by CMB normalization of the scalar power spectrum, $P_{\zeta}(k_{\rm CMB})\approx 2.1\times 10^{-9}$. The squared sound speeds of tensor and scalar modes are positive in both cases of Figure 2. We find taht the scalar perturbations do not experience any amplification in our models and do not contribute to either PBH abundance or secondary GWs.
@@
Let us comment on the amplified frequencies of GWs. Since the sound speed oscillations last for a time period which is much shorter than the duration of inflation, we can identify the characteristic frequency of the amplified GWs with the time when the first stage of inflation ends. This is when the (largest) oscillations of the sound speed take place. The frequency is related to the wavenumber as $k=2\pi f$, while $k$ can be related to the elapsed number of efolds as $k=aH=\beta He^{N}$ (at horizon crossing), where $\beta$ is a constant of normalization of $a$. If we take $N$ as the number of efolds from the horizon exit of the CMB scale $k_{*}$ (i.e. from zero to $\sim 55$), then from $k_{*}=\beta H_{*}$ we can find $\beta$. This allows us to express the frequency as
$f=\frac{k_{*}H}{2\pi H_{*}}e^{N}~{}\Longrightarrow~{}~{}f_{A}=\frac{k_{*}H_{A}}{2\pi H_{*}}e^{N_{A}}~{},$ (43)
where $f_{A}$ is the amplified frequency, while $H_{A}$ and $N_{A}$ are the values when SSR takes place, which can be taken as the end of the first slow-roll phase driven by $\phi$. In the model (42) we have $H_{*}\approx M/\sqrt{6}$, and for an order-of-magnitude estimate of $H_{A}$, we can approximate it by setting $\phi=\dot{\phi}=0$, so that $H_{A}\sim M\gamma\chi_{0}/\sqrt{6}$, where $\chi_{0}$ is the initial value of $\chi$ at which it is frozen until the end of the first slow-roll.
Let us apply Eq. (43) to the parameter choice of Figure 2: $\gamma=0.01$ and $\chi_{0}=9$. In this case, the first slow-roll phase ends at $N\sim 33$, which can be used as the value of $N_{A}$. In Planck units ($c=\hbar=M_{P}=1$) we have ${\rm Mpc}\sim 4\times 10^{56}$ so that $k_{*}=0.05~{}{\rm Mpc}^{-1}\sim 1.2\times 10^{-58}$. Once $f_{A}$ is calculated in Planck units, the Hertz unit can be restored by using its value in Planck units, ${\rm Hz}=1/{\rm s}\sim 2.6\times 10^{-43}$. The result is $f_{A}\sim 0.2~{}{\rm Hz}$, which is a good estimate for the numerical results of Figure 2. Notice that this frequency estimate does not depend on the Gauss–Bonnet parameters, but on the time of transition from the first to the second slow-roll stage. However, the Gauss–Bonnet parameters determine the peak amplitude of the GWs. Although analytical treatment of the GW SSR effect is difficult in the Gauss–Bonnet-coupled inflation, some insights can be drawn from using analytical modeling of the sound speed function, as is done in Refs. [13, 14].
In principle, by testing different scalar potentials and GB functions, one can obtain a larger amplification (in fact, slightly increasing $\gamma$ in (42) can also lead to somewhat larger peaks, as will be shown below) than that of Figure 2. For example, we find that for
@@ VII Conclusions
First, we have demonstrated a time-variation of the sound speeds of scalar and tensor modes driven by the Lagrangian interactions, by focusing on the regime of damped oscillations of the inflaton around its minimum. These oscillations lead to damped oscillations of the tensor mode sound speed, while the variations of the scalar mode sound speed are suppressed by comparison. Then, we showed how the GB-induced SSR can generate a large stochastic background of GWs if the oscillatory phase is followed by more e-folds, as can be realized in two-field double inflation scenario. Notably, we obtained several examples of GW signals which can be tested in future experiments in radio astronomy and space-based interferometers  [^3] An alternative GW phenomenology of (single) GB-coupled inflation, in the frontier of ultra-high-frequencies (UHF) far beyond LIGO, was recently studied in Refs. [58, 59]. Moreover UHF GWs can also have an important impact in radio astronomy from stochastic resonant graviton-photon transitions with cosmological magnetic fields [60, 61].. This represents an important step towards the embedding of the SSR phenomenon, for GWs and PBHs co-genesis, in a theory of inflation based on a Lagrangian principle and fundamental symmetries. Several results show how inflationary $\alpha$-attractors can be obtained from models based on superconformal groups, supergravity and string theory [62, 63, 64], while Gauss–Bonnet portals can be generated by string theory corrections [65, 66, 67, 66]. These encourage the future exploration of SSR in string-inspired inflation scenarios.